Fresh Water, Inc. sold an issue of 6-year $1,000 par value bonds to the public. The bonds have a 8.86 percent coupon rate and pay interest annually. The current market rate of interest on the Fresh Water, Inc. bonds is 12.08 percent. What is the current market price of the bonds
Answer #1

Coupon = 0.0886 * 1,000 = 88.6

Current market price = Coupon * [1 - 1 / (1 + r)n] / r + FV / (1 + r)n

Current market price = 88.6 * [1 - 1 / (1 + 0.1208)6] / 0.1208 + 1000 / (1 + 0.1208)6

Current market price = 88.6 * 4.102109 + 504.465263

Current market price = $867.91

You can also find this using a financial calculator:

FV 1000

PMT 88.6

N 6

I/Y 12.08

CPT PV